OpenStudy (unikornkiller):

Write the decimal expansion for 1/6

OpenStudy (anonymous):

divide 1 by 6

OpenStudy (unikornkiller):

;)

OpenStudy (michele_laino):

please, note that: \[\frac{ 1 }{ 6 }=\frac{ 10/6 }{ 10 }=\frac{ 1+4/6 }{ 10 }=\frac{ 1+2/3 }{ 10 }=\] \[=\frac{ 1 }{ 10 }+\frac{ 2 }{ 3 }*\frac{ 1 }{ 10 }=0.1+0.6*0.1=0.16\]

OpenStudy (unikornkiller):

Is 1.5 a terminating, repeating, or non-repeating decimal?

OpenStudy (michele_laino):

decimals are infinity, namely 0.166666666666.....